# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

### v3

@@ 4 Transformation Functions: Definition and Examples
This section presents appropriate transformations $\boldsymbol{\xi}_{x}=t(x,\boldsymbol{\vartheta})$ with $x\in\mathcal{X}$ for classical distributions in which both Assumptions 3 and 3 hold. We classify the transformation approaches according to four main categories to organize our discussion: enumeration (Section 4.1), inversion (Section 4.2), standardization (Section 4.3), and convolution (Section 4.4). Then, Section 4.5 discusses the proposed methodology’s advantages and disadvantages. We use the notation introduced in Section 2.1, and we denote $\boldsymbol{\xi}_{x}^{u}$ as the $u$-th element of the random vector $\boldsymbol{\xi}_{x}$, and $\boldsymbol{\vartheta}_{u}$ as the random variable or vector used to obtain $\boldsymbol{\xi}_{x}^{u}$. Similarly, $\xi_{x}^{u}$ and $\vartheta_{u}$ denote their realizations. Unless stated otherwise, we consider the random elements of $\boldsymbol{\xi}_{x}$ to be independent. If the random elements of $\boldsymbol{\xi}_{x}$ are independent, then the $\boldsymbol{\vartheta}_{u}$ for $u\in\mathcal{U}$ must also be independent. Thus, we can model each random element $\boldsymbol{\xi}_{x}^{u}$ separately by using its marginal probability distribution conditioned on $x$ and the $\boldsymbol{\vartheta}_{u}$. Nonetheless, in certain cases, when the elements of $\boldsymbol{\xi}_{x}$ are correlated, we can have correlated $\boldsymbol{\vartheta}_{u}$ and also model each $\boldsymbol{\xi}_{x}^{u}$ separately. As a result, we mainly focus on the individual transformation of an element $u$ in this section, i.e., $\boldsymbol{\xi}_{x}^{u}=t(x,\boldsymbol{\vartheta}_{u})$. We remark that most studies have focused on the independent case (see Section (2.2)), while our approach is suitable for both independent and dependent cases.
4.1 Enumeration
For a random element $\boldsymbol{\xi}_{x}^{u}$ with $u\in\mathcal{U}$, a simple transformation can be employed if the set of alternative marginal distributions $\mathcal{D}_{u}$ of element $u$ is sufficiently small to allow enumeration. This corresponds to the “discrete selection of distributions” (see Section 2.1). We have $\boldsymbol{\vartheta}_{u}=\{\boldsymbol{\vartheta}_{ud}\}_{d\in\mathcal{D}_{u}}$, where $\boldsymbol{\vartheta}_{ud}$ is a random variable following distribution $d\in\mathcal{D}_{u}$. The $\boldsymbol{\vartheta}_{u}$ is independent of $x$ and respects Property 1 of Assumption 3. We thus have $t(x,\boldsymbol{\vartheta}_{u})=\sum_{d=1}^{D_{u}}\boldsymbol{\vartheta}_{ud}x_{u}^{d}$, with $x_{u}^{d}$ as defined in Section 2.1, and Property 2 of Assumption 3 holds. The elements of $\boldsymbol{\vartheta}_{u}$ may be independent or correlated (e.g., by using common random numbers). We exemplify this in Example 4.1.
Example 4.1
Consider two independent random elements $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$, with set of realizations $\mathcal{R}_{A}=\{r_{A}^{1},r_{A}^{2},r_{A}^{3}\}$ and $\mathcal{R}_{B}=\{r_{B}^{1},r_{B}^{2},r_{B}^{3}\}$, respectively. Thus, the original program has nine scenarios. These random variables have a set of alternative marginal distributions $\mathcal{D}_{A}=\{d_{A}^{1},d_{A}^{2}\}$ and $\mathcal{D}_{B}=\{d_{B}^{1},d_{B}^{2}\}$, respectively. Only realizations $r_{A}^{1}$ and $r_{A}^{2}$ have a probability greater than zero in distribution $d_{A}^{1}$, while all three realizations have a probability greater than zero in distribution $d_{A}^{2}$ (similarly for $d_{B}^{1}$ and $d_{B}^{2}$). If $d_{A}^{1}$ and $d_{B}^{1}$ are chosen, then only four scenarios have a probability greater than zero. However, if $d_{A}^{2}$ and $d_{B}^{2}$ are chosen, all nine scenarios have positive probabilities. We have two independent exogenous random vectors: $\boldsymbol{\vartheta}_{A}=\{\boldsymbol{\vartheta}_{A,d_{A}^{1}},\boldsymbol{\vartheta}_{A,d_{A}^{2}}\}$ and $\boldsymbol{\vartheta}_{B}=\{\boldsymbol{\vartheta}_{B,d_{B}^{1}},\boldsymbol{\vartheta}_{B,d_{B}^{2}}\}$. While $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ are independent, their elements can be independent or correlated, i.e., $\boldsymbol{\vartheta}_{A,d_{A}^{1}}$ can be correlated with $\boldsymbol{\vartheta}_{A,d_{A}^{2}}$, similarly for $\boldsymbol{\vartheta}_{B,d_{B}^{1}}$ and $\boldsymbol{\vartheta}_{B,d_{B}^{2}}$. The random variable $\boldsymbol{\vartheta}_{A,d_{A}^{1}}$ follows distribution $d_{A}^{1}$ and has two possible realizations ($r_{A}^{1}$ and $r_{A}^{2}$), while $\boldsymbol{\vartheta}_{A,d_{A}^{2}}$ follows distribution $d_{A}^{2}$ and has three possible realizations ($r_{A}^{1}$, $r_{A}^{2}$ and $r_{A}^{3}$). The reasoning is the same for $\boldsymbol{\vartheta}_{B,d_{B}^{1}}$ and $\boldsymbol{\vartheta}_{B,d_{B}^{2}}$. Hence, the transformed program has 36 scenarios. If distributions $d_{A}^{1}$ and $d_{B}^{1}$ are chosen, then $t(x,\boldsymbol{\vartheta}_{A})$ and $t(x,\boldsymbol{\vartheta}_{B})$ return random variables $\boldsymbol{\vartheta}_{A,d_{A}^{1}}$ and $\boldsymbol{\vartheta}_{B,d_{B}^{1}}$, respectively. Thus, for the scenario composed by realizations $r_{A}^{1}$, $r_{A}^{3}$, $r_{B}^{1}$, $r_{B}^{3}$ for random variables $\boldsymbol{\vartheta}_{A,d^{1}_{A}}$, $\boldsymbol{\vartheta}_{A,d^{2}_{A}}$, $\boldsymbol{\vartheta}_{B,d^{1}_{B}}$ and $\boldsymbol{\vartheta}_{B,d^{2}_{B}}$, respectively, the functions $t(x,\boldsymbol{\vartheta}_{A})$ and $t(x,\boldsymbol{\vartheta}_{B})$ return realizations $r_{A}^{1}$ and $r_{B}^{1}$.
If an enumeration of the alternative distributions $\mathcal{D}$ of the entire random vector $\boldsymbol{\xi}$ is viable, then we can apply the same transformation considering all the elements of $\boldsymbol{\xi}$. That is, $t(x,\boldsymbol{\vartheta})=\sum_{d=1}^{D}\boldsymbol{\vartheta}_{d}x^{d}$, where $\boldsymbol{\vartheta}_{d}$ follows distribution $d\in\mathcal{D}$ and $x^{d}$ is defined as in Section 2.1. This transformation can be applied for $\boldsymbol{\xi}_{x}$ with independent or dependent random elements. Finally, if bilinear terms appear due to $t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)}y$, $t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)}x$ and $t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)}y$, then auxiliary linearizations such as McCormick linearization (McCormick 1976) can be applied so Assumption 3 is respected.
@@ 4.2 Inversion
The inversion method is commonly used to sample a random variable when the probability mass or density function is known (e.g., the inverse sampling method, cf. Robert and Casella 1999, Chap. 2). This technique transforms a random variable following the continuous uniform distribution $U(0,1)$ into the desired random variable. We define $\mathbb{F}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ to be the conditional marginal cumulative distribution function of the random variable $\boldsymbol{\xi}^{u}$ given $x$, the first-stage variables, and $\beta$, a set of remaining parameters that are not decisions of the optimization model. To apply the inverse transformation, we compute the generalized inverse of the conditional cumulative distribution:
$\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(j)=\min\left\{\xi\in\Xi_{x}\mid\mathbb{F}_{\boldsymbol{\xi}^{u}}(x,\beta)(\xi)\geq j\right\}.$ (9)
@@
Hence, the transformation is defined as $\boldsymbol{\xi}_{x}^{u}=t(x,\boldsymbol{\vartheta}_{u})=\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\boldsymbol{\vartheta}_{u})$ (Property 2 of Assumption 3), such that $\boldsymbol{\vartheta}_{u}$ is a random variable following $U(0,1)$ and independent of $x$ (Property 1 of Assumption 3). The inversion method, in many cases, has the desirable property of inducing a positive correlation (see Section 3.1) between the estimators of different systems (i.e., the $\hat{g}_{N}(x)$ for $x\in\mathcal{X}$) when their samples are generated from the same sequence of uniform random numbers (Whitt 1976).
@@
If the elements of $\boldsymbol{\xi}_{x}$ are dependent and their correlation matrix is known, then we can define correlated random elements $\boldsymbol{\vartheta}_{u}$ for $u\in\mathcal{U}$ using an appropriate copula. A copula is a multivariate cumulative distribution function that ensures the marginal distribution of each random variable $\boldsymbol{\vartheta}_{u}$ is the $U(0,1)$. Copulas are well-known for modeling dependencies (see, e.g., Nelsen 2006). Thus, as in the independent case, each $\boldsymbol{\vartheta}_{u}$ can be transformed into $\boldsymbol{\xi}_{x}^{u}$ separately using the marginal distribution of this endogenous random element. Example 4.2 illustrates the application of the inverse.
Example 4.2
Consider two independent (correlated) random variables $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$ following the marginal cumulative distribution $\mathbb{F}_{\boldsymbol{\xi}^{A}}(x,\beta)(\cdot)$ and $\mathbb{F}_{\boldsymbol{\xi}^{B}}(x,\beta)(\cdot)$, respectively. We have two independent (correlated) random variables $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ following the marginal distribution $U(0,1)$. For the independent case, the $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ follow the independence bi-variate copula. For the correlated case, the $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ follow a bi-variate copula which models the correlation of $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$, e.g., the Gaussian copula with the corresponding correlation. A scenario is composed of the realizations of $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$, which are transformed into $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$ by their respective transformation $t(x,\boldsymbol{\vartheta}_{u})=\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)$ for $u\in\{A,B\}$.
Next, we investigate the application of this technique for discrete random variables and continuous random variables according to the properties of their cumulative probability distribution.
@@
In this case, the inverse transformation $\boldsymbol{\xi}_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\boldsymbol{\vartheta}_{u})$ is simplified and consists of defining the realization $r\in\mathcal{R}_{u}$ such that
@@
where $\epsilon>0$ is a sufficiently small value used to define the strict inequality in (10). Constraints (11) ensure that the right-hand side in Inequalities (10) is respected. Constraints (12) certify that the left-hand side in Inequalities (10) is respected. Constraints (13) establish the connection between $\pi^{r}_{u}$ and $\theta^{r}_{u}$ so that $\theta^{r}_{u}=1$ for the right realization, where $\pi_{u}^{R_{u}}$ is set to one. Although we could directly model Constraints (11) and (12) with variables $\theta_{u}^{r}$, we use the variables $\pi_{u}^{r}$ to reduce the number of nonzero elements in these constraints. Constraint (14) restricts the choice of a unique realization for the random element $u$. Finally, $\nu_{r}^{u}$ is the value of realization $r$ of the random element $u$. Thus, $t(x,\vartheta_{u})=\sum_{r=1}^{R_{u}}\nu_{r}^{u}\theta_{u}^{r}$. To ensure Assumption 3, the continuous relaxation of Constraints (11) and (12) must be convex. This leads to the following theorem:
@@
If the functions $\hat{p}^{u}_{r}(x)$ for $u\in\mathcal{U}$ and $r\in\mathcal{R}_{u}$ have a concave continuous relaxation on $x\in\mathcal{X}$, Constraints (11) and (12) are convex. The same is true if these functions have a convex continuous relaxation.
Appendix 9 presents the proof of Theorem 4.3. In addition, even if the continuous relaxation of the $\hat{p}_{r}^{u}(x)$ is neither concave nor convex, we can convexify Constraints (11) and (12) in certain cases. This analysis is problem-specific. Different studies investigate problems with convex or concave $p_{r}^{u}(x)$ (see Section 2.2). Despite this desirable property, the original program is usually nonconvex. In contrast, using the inverse transformation function results in only linear or convex constraints.
To the best of our knowledge, only Holzmann and Smith (2021) use the discrete inverse transformation to transform the endogenous uncertainty in 2-SMIPD into exogenous uncertainty for a specific problem. In contrast, we describe a general framework adapted to various functions $\hat{p}_{r}^{u}(x)$. Moreover, the set of Constraints (11)–(14) has a tighter continuous relaxation feasibility set than that used by Holzmann and Smith (2021) (see Appendix 10). Finally, Holzmann and Smith (2021) do not investigate the statistical estimators obtained with their formulation, which we study for general transformations (Section 3) and also validate with a case study in Sections 5 and 6.
In certain cases, such as when $\boldsymbol{\xi}_{x}^{u}$ has only two possible realizations and $p_{r}^{u}(x)$ is a linear function, simpler models can be designed. This is the case of the Bernoulli distribution, described next.
@@
We consider that $\boldsymbol{\xi}_{x}^{u}\sim Bern(\phi)$, with the probability of success $\phi\in[0,1]$ being a subvector of $x$. The inverse is such that $\xi_{x}^{u}=1$ if $\vartheta_{u}<\phi$, and $\xi_{x}^{u}=0$ otherwise, with $\boldsymbol{\vartheta}_{u}\sim U(0,1)$. We model this with the variables $\upsilon_{u}\in\{0,1\}$, i.e., $t(x,\vartheta_{u})=\upsilon_{u}$, and the constraints
@@
For both transformations (general and specific to Bernoulli), linearization techniques, such as McCormick, can be easily applied to linearize the bilinear terms in $t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)}y$, $t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)}x$, and $t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)}y$, as they involve at least one binary variable. Thus, Assumption 3 holds.
@@
In general, the transformation $\xi_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\vartheta_{u})$ cannot be simplified for continuous random variables. Nevertheless, when, for fixed $x$ and $\beta$, the conditional cumulative distribution is strictly increasing and continuous, the generalized inverse is redundant because $\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ is the unique inverse function. This property is commonly respected for well-behaved distributions and leads to a simpler transformation. Two applications of the inverse for some well-known continuous distributions respecting this property are listed below.
@@
If $\boldsymbol{\xi}^{u}\sim U(a,b)$ and both parameters $a\in\mathbb{R}$ and $b\in\mathbb{R}$ are first-stage variables, then we can use the transformation function $t(a,b,\boldsymbol{\vartheta}_{u})=a+(b-a)\boldsymbol{\vartheta}_{u}$.
@@
If $\boldsymbol{\xi}^{u}\sim\exp(\lambda)$ with $\lambda>0$ a first-stage decision, then the transformation is defined as $t(\lambda,\boldsymbol{\vartheta}_{u})=-{\ln(1-\boldsymbol{\vartheta}_{u})}/{\lambda}$. This function is concave for $\lambda>0$.
Both distributions are part of the special case where the unique inverse $\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ is convex or concave on $x\in\mathcal{X}$, i.e., the inverse function is convex or concave on the parameters of the distribution. Other distributions that fall into this special case are the Cauchy and logistic distributions. Additionally, there are distributions for which $\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ has no closed-form expression, but good numerical approximations are available, e.g. student, chi-square, and normal distributions. Nevertheless, transformations based on other techniques, which lead to simpler formulations, can be used for certain distributions (see Section 4.3 for the normal distribution). The analysis of the special class of distributions introduced before leads to the following theorem:
@@
Assumption 3 holds if the following criteria are met: (1) the cost vector $q$, technology matrix $T$ and recourse matrix $W$ are fixed; (2) $\mathbb{F}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$for fixed $x$ and $\beta$ is continuous and strictly increasing; and (3) the transformation $t(x,\boldsymbol{\vartheta}_{u})=\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ is a convex function on $x\in\mathcal{X}$, or concave if the recourse constraint in Program (6b) is an inferior or equal to inequality.
Appendix 9 presents the proof of Theorem 4.4. If Criterion (1) does not hold, the products $t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)}y$, $t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)}x$ and $t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)}y$ may be nonconvex even if Criterion (3) holds. Thus, a problem-specific convexity analysis is required to confirm if Assumption 3 is satisfied.

### v4

@@ 4 Transformation Functions: Definition and Examples
This section presents appropriate transformations $\bm{\xi}_{x}=t(x,\bm{\vartheta})$ with $x\in\mathcal{X}$ for classical distributions in which both Assumptions 3 and 3 hold. We classify the transformation approaches according to four main categories to organize our discussion: inversion (Section 4.1), enumeration (Section 4.2), standardization (Section 4.3), and convolution (Section 4.4). Then, Section 4.5 discusses the proposed methodology’s advantages and disadvantages. We use the notation introduced in Section 2.1, and we denote $\bm{\xi}_{x}^{u}$ as the $u$-th element of the random vector $\bm{\xi}_{x}$, and $\bm{\vartheta}_{u}$ as the random variable or vector used to obtain $\bm{\xi}_{x}^{u}$. Similarly, $\xi_{x}^{u}$ and $\vartheta_{u}$ denote their realizations. Unless stated otherwise, we consider the random elements of $\bm{\xi}_{x}$ to be independent. If the random elements of $\bm{\xi}_{x}$ are independent, then the $\bm{\vartheta}_{u}$ for $u\in\mathcal{U}$ must also be independent. Thus, we can model each random element $\bm{\xi}_{x}^{u}$ separately by using its marginal probability distribution conditioned on $x$ and the $\bm{\vartheta}_{u}$. Nonetheless, in certain cases, when the elements of $\bm{\xi}_{x}$ are correlated, we can have correlated $\bm{\vartheta}_{u}$ and also model each $\bm{\xi}_{x}^{u}$ separately in the same way as in the independent case, as explained further in this section. As a result, we mainly focus on the individual transformation of an element $u$ in this section, i.e., $\bm{\xi}_{x}^{u}=t(x,\bm{\vartheta}_{u})$. We remark that most studies have focused on the independent case (see Section (2.2)), while our approach is suitable for both independent and dependent cases.
@@ 4.1 Inversion
The inversion method is commonly used to sample a random variable when the probability mass or density function is known (e.g., the inverse sampling method, cf. Robert and Casella 1999, Chap. 2). This technique transforms a random variable following the continuous uniform distribution $U(0,1)$ into the desired random variable. We define $\mathbb{F}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ to be the conditional marginal cumulative distribution function of the random variable $\bm{\xi}^{u}$ given $x$, the first-stage variables, and $\beta$, a set of remaining parameters that are not decisions of the optimization model. To apply the inverse transformation, we compute the generalized inverse of the conditional cumulative distribution:
$\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(j)=\min\left\{\xi\in\Xi_{x}\mid\mathbb{F}_{\bm{\xi}^{u}}(x,\beta)(\xi)\geq j\right\}.$ (8)
@@
The transformation is then defined as $\bm{\xi}_{x}^{u}=t(x,\bm{\vartheta}_{u})=\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\bm{\vartheta}_{u})$ (Property 2 of Assumption 3), such that $\bm{\vartheta}_{u}$ follows the $U(0,1)$ and is independent of $x$ (Property 1 of Assumption 3). The inversion method, in many cases, has the desirable property of inducing a positive correlation (see Section 3.1) between the estimators of different systems (i.e., the $\hat{g}_{N}(x)$ for $x\in\mathcal{X}$) when their samples are generated from the same sequence of uniform random numbers (Whitt 1976).
@@
If the elements of $\bm{\xi}_{x}$ are dependent and their correlation matrix is known, then we can define correlated random elements $\bm{\vartheta}_{u}$ for $u\in\mathcal{U}$ using an appropriate copula. A copula is a multivariate cumulative distribution function that ensures the marginal distribution of each random variable $\bm{\vartheta}_{u}$ is the $U(0,1)$. Copulas are well-known for modeling dependencies (see, e.g., Nelsen 2006). Thus, as in the independent case, each $\bm{\vartheta}_{u}$ can be transformed into $\bm{\xi}_{x}^{u}$ separately using the marginal distribution of this endogenous random element (i.e., the transformation is equal for both independent and dependent cases; the difference lies in the definition of $\bm{\vartheta}_{u}$). To keep the paper self-contained, we illustrate the application of the inverse in Appendix 9 (Example 9.1) for both cases. We note that this method is limited to cases where the function $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(j)$ has a well-defined or approximated form. Next, we investigate the application of this technique for discrete random variables and continuous random variables according to the properties of their cumulative probability distribution.
@@
In this case, the inverse transformation $\bm{\xi}_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\bm{\vartheta}_{u})$ is simplified and consists of defining the realization $r\in\mathcal{R}_{u}$ such that
@@
where $\epsilon>0$ is a sufficiently small value used to define the strict inequality in (9). Constraints (10) and (11) ensure that the right- and left-hand side in Inequalities (9) are respected, respectively. Constraints (12) establish the connection between $\pi^{r}_{u}$ and $\theta^{r}_{u}$ so that $\theta^{r}_{u}=1$ for the right realization, where $\pi_{u}^{R_{u}}$ is set to one. Although we could model Constraints (10) and (11) with variables $\theta_{u}^{r}$, we use the variables $\pi_{u}^{r}$ to reduce the number of nonzero elements in these constraints. Constraint (13) restricts the choice of a unique realization for element $u$. Finally, $\nu_{r}^{u}$ is the value of realization $r$ of the random element $u$, so $t(x,\vartheta_{u})=\sum_{r=1}^{R_{u}}\nu_{r}^{u}\theta_{u}^{r}$. To ensure Assumption 3, the continuous relaxation of Constraints (10) and (11) must be convex. This leads to the following theorem:
@@
If all the functions $\hat{p}^{u}_{r}(x)$ for $u\in\mathcal{U}$ and $r\in\mathcal{R}_{u}$ are either convex or concave over $\left[0,1\right]^{n_{1}}$, then Constraints (10) and (11) have a convex continuous relaxation.
The Appendix 8 presents the proof of Theorem 4.1. Different studies investigate problems with convex or concave $p_{r}^{u}(x)$ (see Section 2.2). Despite this desirable property, the original program is usually nonconvex. In contrast, using the inverse transformation function results in only linear or convex constraints. In addition, even if the continuous relaxation of the $\hat{p}_{r}^{u}(x)$ is neither concave nor convex, we can convexify Constraints (10) and (11) in certain cases (this analysis is problem-specific). Nonetheless, in addition to relying on the probability function $\hat{p}^{u}_{r}(x)$ to ensure convexity, the number of integer variables and constraints increases polynomially with the number of random parameters $U$ and their realizations $R_{u}$, which impacts the problem’s complexity and highlights the computational challenges of handling endogenous uncertainty.
To the best of our knowledge, only Holzmann and Smith (2021) used inversion to model a specific discrete endogenous random variable in 2-SMIPD. In contrast, we describe a general framework adapted to various functions $\hat{p}_{r}^{u}(x)$. The set of Constraints (10)–(13) also has a tighter continuous relaxation feasibility set than that employed by Holzmann and Smith (2021) (see Appendix 10).
In certain cases, such as when $\bm{\xi}_{x}^{u}$ has only two possible realizations and $p_{r}^{u}(x)$ is a linear function, simpler models can be designed. This is the case of the Bernoulli distribution, described next.
@@
We consider that $\bm{\xi}_{x}^{u}\sim Bern(\phi)$, with the probability of success $\phi\in[0,1]$ being a subvector of $x$. The inverse is such that $\xi_{x}^{u}=1$ if $\vartheta_{u}<\phi$, and $\xi_{x}^{u}=0$ otherwise, with $\bm{\vartheta}_{u}\sim U(0,1)$. We model this with the variables $\upsilon_{u}\in\{0,1\}$, i.e., $t(x,\vartheta_{u})=\upsilon_{u}$, and the constraints
@@
For both transformations (general and specific to Bernoulli), if only $h(\omega)$ is uncertain, then Assumption 3 is respected. Otherwise, linearization techniques, such as McCormick, can be applied to linearize the bilinear terms in $t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)}y$, $t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)}x$, and $t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)}y$, so Assumption 3 holds. In this case, the problem’s complexity increases further, and the transformation is restricted by the dimensions of variables $x$ and $y$, due to the inclusion of extra variables and constraints when applying linearization techniques.
@@
In general, the transformation $\xi_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\vartheta_{u})$ cannot be simplified for continuous random variables. Nevertheless, when, for fixed $x$ and $\beta$, the conditional cumulative distribution is strictly increasing and continuous, the generalized inverse is redundant because $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ is the unique inverse function. This property is commonly respected for well-behaved distributions and leads to a simpler transformation. Two applications of the inverse for some well-known continuous distributions respecting this property are listed below.
@@
If $\bm{\xi}^{u}\sim U(a,b)$ and both parameters $a\in\mathbb{R}$ and $b\in\mathbb{R}$ are first-stage variables, then we can use the transformation function $t(a,b,\bm{\vartheta}_{u})=a+(b-a)\bm{\vartheta}_{u}$.
@@
If $\bm{\xi}^{u}\sim\exp(\lambda)$ with $\lambda>0$ a first-stage decision, then the transformation is defined as $t(\lambda,\bm{\vartheta}_{u})=-{\ln(1-\bm{\vartheta}_{u})}/{\lambda}$. This function is concave for $\lambda>0$.
Both distributions are part of the special case where the unique inverse $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ is convex or concave on $x\in\mathcal{X}$, i.e., the inverse function is convex or concave on the parameters of the distribution. Other distributions that fall into this special case are the Cauchy and logistic distributions. Additionally, there are distributions for which $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ has no closed-form expression, but good numerical approximations are available, e.g. student, chi-square, and normal distributions. Nevertheless, transformations based on other techniques, which lead to simpler formulations, can be used for certain distributions (see Section 4.3 for the normal distribution). The analysis of the special class of distributions introduced before leads to the following theorem:
@@
Assumption 3 holds if the following criteria are met: (1) the cost vector $q$, technology matrix $T$ and recourse matrix $W$ are fixed; (2) $\mathbb{F}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$for fixed $x$ and $\beta$ is continuous and strictly increasing; and (3) the transformation $t(x,\bm{\vartheta}_{u})=\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ is a convex function on $x\in\mathcal{X}$, or concave if the recourse constraint in Program (5) is an inferior or equal to inequality.
The Appendix 8 presents the proof of Theorem 4.2. If Criterion (1) does not hold, the products $t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)}y$, $t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)}x$ and $t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)}y$ may be nonconvex even if Criterion (3) holds. Thus, a problem-specific convexity analysis is required to confirm if Assumption 3 is satisfied. This evidences the complexity added by endogenous uncertainty, as the transformation can, in some cases, produce a nonconvex nonlinear transformed program with exogenous uncertainty, which is theoretically intractable.
4.2 Enumeration
For a random element $\bm{\xi}_{x}^{u}$ with $u\in\mathcal{U}$, a simple transformation can be employed if the set of alternative marginal distributions $\mathcal{D}_{u}$ of element $u$ is sufficiently small to allow enumeration. This corresponds to the “discrete selection of distributions” (see Section 2.1). We have $\bm{\vartheta}_{u}=\{\bm{\vartheta}_{ud}\}_{d\in\mathcal{D}_{u}}$, where $\bm{\vartheta}_{ud}$ is a random variable following distribution $d\in\mathcal{D}_{u}$. The $\bm{\vartheta}_{u}$ is independent of $x$ (Property 1 of Assumption 3). We thus have ${t(x,\bm{\vartheta}_{u})=\sum_{d=1}^{D_{u}}\bm{\vartheta}_{ud}x_{u}^{d}}$ (Property 2 of Assumption 3), with $x_{u}^{d}$ as defined in Section 2.1. This transformation can be applied for $\bm{\xi}_{x}$ with independent or correlated random elements by also using copulas. To keep the paper self-contained, we illustrate this in Appendix 9, Example 9.2, and also present the case where the distribution is selected for the entire random vector $\bm{\xi}_{x}$ in Example 9.3. If only $h(\omega)$ is uncertain, then Assumption 3 is directly satisfied. Otherwise, bilinear terms may appear due to $t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)}y$, $t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)}x$ and $t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)}y$. In such cases, auxiliary linearizations such as the McCormick envelope can be applied so Assumption 3 is respected. However, this approach is limited not only by the size of $\mathcal{D}_{u}$ but also by the dimensions of $y$ or $x$, as it increases the number of variables, leading to computational challenges.